OpenStudy (anonymous):

James added 3 and two-thirds plus 4 and one-eighth. His work is shown here. What did James do right? What did he do wrong? What should he have done? three and two-thirds plus 4 and one-eighth equals 7 and three-elevenths

OpenStudy (anonymous):

2/3 + 1/8 = 16/24 + 3/24 = 19/24. 3 + 4 = 7 7 19/24 He added the whole numbers right but added the fractions wrong.